At Isogen Pharmaceuticals, the filling process for its asthma inhaler is set to dispense 150 milliliters (ml) of steroid solution per container. The average range for a sample of 4 containers is 3 ml. The lower control limit for the mean chart, rounded to three decimal places, is ……….. milliliters.
Answer #1

Solution:

Mean (X-bar) = 150 ml

Average range (R-bar) = 3 ml

Sample size (n) = 4 containers

From the table of factors for computing control chart limits (3-sigma),

for n = 4 (sample size = 4),

A2 = 0.729

The Lower Control Limit (LCL) for the mean chart is calculated as,

LCL = X-bar - (A2 x R-bar)

LCL = 150 - (0.729 x 3)

LCL = 147.813 millilitres

> for what is give it is correct but the question is not all the way full.
D4=2.282 from the table given for 4
D3=0 for the table given for 4
R=3
UCL=D4*R=6.85
LCLr=0
n=150
A2=0.729 for four given in the table
UCLn=R*A2+n=3*.729+150=152.19
LCLn=150-3*.729=147.81

Yr Sun, Oct 29, 2023 11:06 AM